# A Generalization of Lévy’s Theorem on Transition Matrices

### 07 Jan 2024 v1

@@ 3. Operator Semigroups on Discrete Spaces
which is an ideal in $\ell^{0}$ and hence itself a discrete Archimedean Riesz space that contains the space $c_{00}$ of all finite sequences. On this space, we define the mappings $T_{t}(x_{n})\coloneqq\big{(}\sum_{n}x_{n}p_{nm}(t)\big{)}_{m}$. By definition of $E$, $T_{t}\colon E\to E$ is an one-parameter semigroup that satisfies all conditions of Corollary 3.1. Hence, for every $n,m\in\mathds{N}$, $p_{nm}(t)=\langle T_{t}e_{n},e_{m}\rangle$ is either strictly positive for all $t>0$ or identically zero.
@@ Appendix A Analytic Semigroups
In what follows, a one-parameter operator semigroup $(T_{t})_{t\in(0,\infty)}$ on a Banach space $X$ is called analytic (or holomorphic) if there exists $\theta\in(0,\frac{\pi}{2}]$ such the operator valued mapping $t\mapsto T_{t}$ has an analytic extension to $\Sigma_{\theta}\coloneqq\{z\in\mathds{C}\setminus\{0\}:\lvert\arg z\rvert<\theta\}$. In the literature the definition of analyticity typically requires in addition a boundedness condition in the neighborhood of the origin, cf. [1, Def 3.7.1] or [12, Ch 2], which is not needed in this section. As the most important example, we mention that on a Hilbert space every semigroup generated by a self-adjoint operator that is bounded above is analytic, cf. [8, Cor II.4.7]. For the sake of completeness, we recall that a semigroup $(T_{t})_{t\in(0,\infty)}$ is called a $C_{0}$-semigroup if $\lim_{t\downarrow 0}T_{t}x=x$ for all $x\in X$. An interesting example of an analytic semigroup that is not $C_{0}$ is for instance given by heat euqation with non-local boundary conditions on $L^{\infty}(\Omega)$, cf. [3, Thm 4.8].
@@
Let $E$ be a Banach lattice and let $(T_{t})_{t\in(0,\infty)}$ be an analytic positive semigroup on $E$. Let $f\in E_{+}$ and $\varphi\in E^{\prime}_{+}$. Then either $\langle\varphi,T_{t}f\rangle=0$ for all $t\in(0,\infty)$ or $\langle\varphi,T_{t}f\rangle>0$ for all $t\in(0,\infty)$.
A result in the same vein was proven in [4, Prop 3.12] for eventually positive rather than positive semigroups; while the semigroups considered there are $C_{0}$, the fact that eventual positivity suffices indicates that the behaviour at time 00 is not important. The proof of Thm A.1 is an adaptation of the classical proof for $C_{0}$-semigroups and was communicated to the author by Jochen Glück.
@@
Let $\langle\varphi,T_{s}f\rangle=0$ for some time $s>0$. Step 1: Assume, in addition, that $T_{t}f\downarrow f$ as $t\downarrow 0$. Then we can choose sequences $(t_{n})$ in $(0,s)$ and $(f_{n})$ in $E_{+}$ as in Lemma A.2. For each fixed index $m$ and all $n\geq m$ we get
@@
Step 1 in the proof above is simply the argument for the $C_{0}$-semigroup case. The main point is to observe that this argument does not really need the entire semigroup to be $C_{0}$, but just the orbit of $f$.

### 17 Mar 2024 v2

@@ 3. Operator Semigroups on Sequence Spaces
is a sublattice of $\ell^{0}$ (even an ideal) that contains all unit vectors. On this space, we define the mappings $T_{t}x\coloneqq\big{(}\sum_{n}x_{n}p_{nm}(t)\big{)}_{m\in\mathds{N}}$. Using the definition of $E$ and the properties of the positive matrix semigroup $p_{nm}$, it is easy to check that $T_{t}\colon E\to E$ defines a one-parameter semigroup that satisfies all conditions of Corollary 3.1. Hence, for every $n,m\in\mathds{N}$, $p_{nm}(t)=(T_{t}e^{n})_{m}$ is either strictly positive for all $t>0$ or identically zero.
A norm on a vector lattice satisfying $\lVert x\rVert\leq\lVert y\rVert$ whenever $\lvert x\rvert\leq\lvert y\rvert$ is called a lattice norm; a vector lattice endowed with a complete lattice norm is called a Banach lattice. A lattice norm on a sequence space is called order continuous if every order convergent sequence converges in norm. For instance, $c_{0}$ and $\ell^{p}$ for $1\leq p<\infty$ are sequence spaces with complete and order continuous lattice norms. On the other hand, the sup norm on $\ell^{\infty}$ is a complete lattice norm which is not order continuous.
On a sequence space endowed with a complete and order continuous lattice norm, the continuity conditions on the semigroup in Corollary 3.1 can be formally weakened: as in the case for Markovian matrix semigroups, the measurability of the components already implies their continuity.
Corollary 3.3.
Let $E$ be a sequence space with a complete and order continuous lattice norm and let $(T_{t})_{t\in(0,\infty)}$ be a positive semigroup on $E$. If $t\mapsto(T_{t}e^{i})_{j}$ is measurable on $(0,\infty)$ for all $i,j\in\mathds{N}$, then the following assertions hold:
  1. (a)
    For all $x\in E$ the mapping $t\mapsto T_{t}x$ is continuous on $(0,\infty)$.
  2. (b)
    For all $i,j\in\mathds{N}$ either $(T_{t}e^{i})_{j}>0$ for all $t>0$ or $(T_{t}e^{i})_{j}=0$ for all $t>0$.
Proof.
As every positive operator on a complete normed vector lattice is continuous, cf. [19, Prop 1.3.5], and the norm is order continuous, every $T_{t}$ is order continuous.
Let $x\in E_{+}$. It follows from the order continuity of all $T_{t}$ that
$t\mapsto(T_{t}x)_{j}=\sup_{N}\sum_{k=1}^{N}x_{k}(T_{t}e^{k})_{j}$
is measurable on $(0,\infty)$ for all $j\in\mathds{N}$. Now let $\varphi\in E^{\prime}$ be a positive functional. Due to the order continuity of the norm, $\varphi$ is order continuous and therefore
$t\mapsto\langle\varphi,T_{t}x\rangle=\sup_{N}\sum_{k=1}^{N}(T_{t}x)_{k}\langle\varphi,e^{k}\rangle$
is measurable on $(0,\infty)$. Decomposing general $x\in E$ and $\varphi\in E^{\prime}$ in positive and negative parts shows that $t\mapsto T_{t}$ is weakly measurable. As $E$ is separable, $t\mapsto T_{t}$ is strongly measurable. Now assertion (a) follows from [11, Thm 10.2.3].
Assertion (b) now is an immediate consequence of (a) and Corollary 3.1. ∎
Remark 3.4.
The proof of Part (a) of Corollary 3.3 also shows that a positive matrix semigroup $p_{ij}$ which is merely known to be measurable on $(0,\infty)$ is automatically continuous on $(0,\infty)$ if its associated one-parameter semigroup as constructed in Remark 3.2 preserves some sequence space with a complete and order continuous norm such as $c_{0}$ or $\ell^{p}$ for $1\leq p<\infty$. This observation also includes the well-known fact that every Markovian matrix semigroup is continuous, cf. [8, Thm II.1].
While the proof of Theorem 2.4 makes heavy use of the fact that the index set is at most countable, the statement holds true even on more general discrete spaces with possibly uncountably many atoms whenever the orbits of the considered semigroup necessarily vanishes on all but countably many atoms. We show in the following corollary that this is the case for positive and strongly continuous one-parameter semigroups on one of the following spaces: $\ell^{p}(\mathds{R})$ for some $1\leq p<\infty$ or
$c_{0}(\mathds{R})\coloneqq\{f\colon\mathds{R}\to\mathds{R}:\text{ for all }\varepsilon>0\text{ the set }\{r\in\mathds{R}:\lvert f(r)\rvert>\varepsilon\}\text{ is finite}\}$
endowed with the sup norm. On both spaces the proof is based on the fact that
$\inf_{F\subseteq\mathds{R}\text{ finite}}\lVert f\cdot\mathds{1}_{\mathds{R}\setminus F}\rVert=0$
for all vectors $f$.
Corollary 3.5.
Let either $E=\ell^{p}(\mathds{R})$ for some $1\leq p<\infty$ or $E=c_{0}(\mathds{R})$ and let $(T_{t})_{t\in(0,\infty)}$ be a positive semigroup on $E$ such that $t\mapsto T_{t}f$ is continuous on $(0,\infty)$ for all $f\in E$. Then the following assertions hold:
  1. (a)
    For each $f\in E$ the set $\{r\in\mathds{R}:\lvert T_{t}f\rvert(r)>0\text{ for some }t>0\}$ is at most countable.
  2. (b)
    For all $r,s\in\mathds{R}$ either $(T_{t}e^{r})(s)>0$ for all $t>0$ or $(T_{t}e^{r})(s)=0$ for all $t>0$. As before, $e^{r}=\mathds{1}_{\{r\}}$ denotes the indicator function of the singleton $\{r\}$.
Proof.
(a) Let $f\in E$ and let $K\subseteq(0,\infty)$ be a compact interval. Let $\delta\geq 0$. We are going to show that the set
$R_{\delta}\coloneqq\{r\in\mathds{R}:\lvert T_{t}f\rvert(r)>\delta\text{ for some }t\in K\}$
is finite. Using the notation $\mathscr{F}\coloneqq\{F\subseteq\mathds{R}:F\text{ finite}\}$ we have $\inf_{F\in\mathscr{F}}\lVert(T_{t}f)\mathds{1}_{\mathds{R}\setminus F}\rVert=0$ for all $t\in K$. Moreover, since for all $F\in\mathscr{F}$, $\lVert(T_{t}f)\mathds{1}_{\mathds{R}\setminus F}\rVert$ is continuous in $t$, the sets
$K_{F}\coloneqq\{t\in K:\lVert(T_{t}f)\mathds{1}_{\mathds{R}\setminus F}\rVert<\delta/2\}$
form an open cover of $K$. By compactness of $K$ and because $K_{F}$ increases with $F$, we find $F\in\mathscr{F}$ such that $K\subseteq K_{F}$. In particular, $\lvert T_{t}f\rvert(r)<\delta/2$ for all $r\in\mathds{R}\setminus F$ and $t\in K$. This implies that $R_{\delta}\subseteq F$ showing that $R_{\delta}$ is finite. Since $\delta>0$ and $K\subseteq(0,\infty)$ were arbitrary, this implies assertion (a).
(b) Let $r,s\in\mathds{R}$. By Part (a), $R\coloneqq\{v\in\mathds{R}:(T_{t}e^{r})(v)>0\text{ for some }t>0\}$ is at most countable. As each $T_{t}$ is order continuous, the subspace $\ell^{p}(R)=\{f\in\ell^{p}(\mathds{R}):f(v)=0\text{ for all }v\in\mathds{R}\setminus R\}$ is invariant under each operator $T_{t}$. Hence, for $s\in R$ the assertion follows from Part (b) of Corollary 3.3 applied to the restriction of $(T_{t})_{t\in(0,\infty)}$ to $\ell^{p}(R)$; for $r\in\mathds{R}\setminus R$ one trivially has $(T_{t}e^{r})(s)=0$ for all $t>0$. ∎
@@ Appendix A Analytic Semigroups
In what follows, a one-parameter operator semigroup $(T_{t})_{t\in(0,\infty)}$ on a Banach space $X$ is called analytic (or holomorphic) if there exists $\theta\in(0,\frac{\pi}{2}]$ such the operator valued mapping $t\mapsto T_{t}$ has an analytic extension to $\Sigma_{\theta}\coloneqq\{z\in\mathds{C}\setminus\{0\}:\lvert\arg z\rvert<\theta\}$. It is worth mentioning that the mapping $\Sigma_{\theta}\ni z\mapsto T_{z}$ is analytic if and only if $\Sigma_{\theta}\ni z\mapsto\langle T_{z}x,\varphi\rangle$ is analytic for all $x\in X$ and $\varphi\in X^{\prime}$. In the literature the definition of analyticity typically requires in addition a boundedness condition in the neighborhood of the origin, cf. [1, Def 3.7.1] or [14, Ch 2], which is not needed in this section. As the most important example, we mention that on a Hilbert space every semigroup generated by a self-adjoint operator that is bounded above is analytic, cf. [9, Cor II.4.7]; as the most important counterexample we mention that the shift semigroup on $L^{p}(\mathds{R})$ is not analytic for any $1\leq p<\infty$ (and Theorem A.1 fails to hold). For the sake of completeness, we recall that a semigroup $(T_{t})_{t\in(0,\infty)}$ is called a $C_{0}$-semigroup if $\lim_{t\downarrow 0}T_{t}x=x$ for all $x\in X$. An interesting example of an analytic semigroup that is not $C_{0}$ is for instance given by heat equation with non-local boundary conditions on $L^{\infty}(\Omega)$, cf. [3, Thm 4.8].
@@
Let $E$ be a Banach lattice and let $(T_{t})_{t\in(0,\infty)}$ be a positive and analytic semigroup on $E$. Let $f\in E_{+}$ and $\varphi\in E^{\prime}_{+}$. Then either $\langle\varphi,T_{t}f\rangle=0$ for all $t\in(0,\infty)$ or $\langle\varphi,T_{t}f\rangle>0$ for all $t\in(0,\infty)$.
A result in the same vein was proven in [4, Prop 3.12] for eventually positive rather than positive semigroups; while the semigroups considered there are $C_{0}$, the fact that eventual positivity suffices indicates that the behavior at time 00 is not important. The proof of Thm A.1 is an adaptation of the classical proof for $C_{0}$-semigroups and was communicated to the author by Jochen Glück.
@@
Let $\langle\varphi,T_{s}f\rangle=0$ for some time $s>0$. Step 1: Assume, in addition, that $T_{t}f\to f$ as $t\downarrow 0$. Then we can choose sequences $(t_{n})$ in $(0,s)$ and $(f_{n})$ in $E_{+}$ as in Lemma A.2. For each fixed index $m$ and all $n\geq m$ we get
@@
Step 1 in the proof above is simply the argument for the $C_{0}$-semigroup case; the main point is to observe that this argument does not really need the entire semigroup to be $C_{0}$ but just the orbit of $f$.